OpenStudy (anonymous):

I really suck at math pls help find all zeros of the equation -3x^4+27x^2+1200=0

OpenStudy (anonymous):

\[-3 x^4+27 x^2+1200=-3 (x-5) (x+5) \left(x^2+16\right) \]

OpenStudy (anonymous):

is there more to this answer? or does it stop there?